OpenStudy (anonymous):

Solve the equation: |3x-6|=15

OpenStudy (anonymous):

\[|3x-6|=15\] \[\Rightarrow 3x - 6 = 15 \implies 3x = 21 \implies 21 \div 3 = 7\] \[\Rightarrow 3x - 6 = -15 \implies 3x = -9 \implies -9 \div 3 = -3\] So: \[x = -3,~7\] http://www.regentsprep.org/Regents/math/algtrig/ATE1/abslesson.htm